Currently stuck on preparing a classified balance sheet

Required information The following information applies to the questions displayed below] The following accounts are taken from the financial statements of Facebook Inc. at September 30, 2016. (Amounts f 3 are in millions.) Accounts Payable Accounts Receivable Cash Common Stock Equipment Income Tax Expense Interest Expense Notes Payable (long-term) Prepaid Rent Retained Earnings Service Revenue Short-Term Investments Software 1,950 ok 4,1e0 14.480 4,150 25e 1,988 2,100 4,600 15,3e0 8,388 3,700 < Prev 11 12 13 | of 16 Next >
Part 1 of 3 At Se 30, 2016 0 56 points otal Noncurrent Assets Assets くPrex 111213|of16 Next >
Answer #1
BALANCE SHEET
in $ million
Current Assets: Current Liabilities
a Cash $4,100 Accounts Payable $1,950 j
b Accounts Receivable $500 Total Current Liabilities $1,950 k
c Prepaid Rent $2,100
d Short term investments $8,300
e=a+b+c+d Total Current Assets $15,000 Long term Liabilities:
Non current Assets Notes payable- Long Term $1,900 l
f Equipment $4,150 Total Long term Liabilities $1,900 m
g Software $3,700
h=f+g Total Non current Assets $7,850 Total Liabilities $3,850 n=k+m
Stockhoders' Equity
Common Stock $14,400 p
Retained Earnings $4,600 q
Total Stockholders' Equity $19,000 r=p+q
i=e+h TOTAL ASSETS $22,850 Total Liabilities & Stockholders' Equity $22,850 s=n+r